A,B,C are three points in a straight horizontal line.
A plane is flying horizontally directly above ABC.
When the plane passes over point B, the angle of elevation from A is 42 degrees.
When the plane passes over point C, The angle of elevation from A is 34 degrees.
The horizontal distance between B and C is 2km.
Find the height of the plane above ABC to the nearest kilometre.

Answer by Boreal(15235)

==*====*=====*
==A====B====C
tan 42=h/x
tan 34=h/(x+2)
0.900=h/x
h=0.900x
(x+2)*0.675=h
set the two equal
0.900x=0.675x+1.350
0.225x=1.350
x=6 km
Therefore h=0.900x=5.4 km